OpenStudy (anonymous):

what is the molecular formula of C3H2N with 260.24g/mol

OpenStudy (anonymous):

i can just tell that it is C15H10N5 cant tell you the exact molecular formula

OpenStudy (jfraser):

\(C_3H_2N\) tells you the lowest possible whole number ratio of atoms in that molecule. That EMPIRICAL FORMULA must have the same ratio as the actual, MOLECULAR FORMULA. The empirical mass and molecular mass must share the same ratio as the formulas do, so find the mass of \(C_3H_2N\): \(3(12g) + 2(1g) + 14g = 52g\) and compare that to the molecular weight of 260g. \[\frac{MF}{EF} = \frac{MW}{EW}\]so\[\frac{MF}{EF} = \frac{260g}{52g} = 5\]the actual formula is five times heavier than the empirical formula, so take the empirical formula and multiply each subscript by 5. That's why it works

OpenStudy (anonymous):

@JFraser can we tell the name of molecule from above data

OpenStudy (jfraser):

@mekaprasantho not necessarily. with large organic molecules, there are several ways to assemble that many atoms into different molecules. If we want a name, we have to know something else about the molecule. We'd need some bond information, like an IR spectrum, or NMR data, possibly.